OpenStudy (anonymous):

Simplify the expression. −4y + 7y − 5y + 14y =

OpenStudy (math&ing001):

Add them to each other.

OpenStudy (anonymous):

f -3y2-19y2=-21y2?

OpenStudy (anonymous):

so -21y2?

OpenStudy (math&ing001):

Nope -4y + 7y - 5y + 14y = (-4+7-5+14)y

OpenStudy (anonymous):

so if you add whats inside those () it would be..-3-19...

OpenStudy (math&ing001):

Not really, don't you know how to add numbers ?

OpenStudy (math&ing001):

-4+7-5+14 = 12